One student is pushing on a chair with a force of 240.8 N directed at an angle of 30 degrees above horizontal while a second student pushes on the same side of the chair with a force of 107.1 N at an angle of 15 degrees below horizontal. What is the magnitude, in N, of the sum of these forces?
Answer #1

net force is given by

F^2 = F1^2 + F2^2 + 2 F1 F2 cos x

F^2 = 240.8^2 + 107.1^2 + 2* 240.8^2 + 107.1* cos ( 30 + 15)

F = 325.46 N

=========

Comment before rate in case any doubt, will reply for sure.. goodluck